OpenStudy (anonymous):

What is the equation of the line in slope-intercept form that passes through the point (-2, 4) and is perpendicular to the line y = -2x + 4?

OpenStudy (zzr0ck3r):

we know if we have slope m then a line perpendicular will have -1/m slope so your slope will be 1/2 y=(1/2)x+b now use that point to find b 4=(1/2)(-2)+b 4=-1+b b=5 so y=(1/2)x+5

OpenStudy (zzr0ck3r):

or maybe I gave you the completely wrong answer to make sure your paying attention...and not just trying to find answers:)

OpenStudy (anonymous):

hopefully not because i have like 3 more questions and i have to submit this assignment before 12 i have 8 minutes!!!

OpenStudy (zzr0ck3r):

we are not here to do the work for you, we are here to show you how to do it yourself.

OpenStudy (zzr0ck3r):

you seem to pretty much paste the same question over and over, are you learning anything from the answers?

OpenStudy (anonymous):

i understand that but in a time like this.. i just need to pass! but i do appriciate it when you explain it i actually understand it! thanks.